# Universidad de Granada, Campus de Fuentenueva, E-18071 Granada, Spain

### v1

@@ 3 Scalar Schwinger current in de Sitter
$\begin{split}S=\int d^{4}x\sqrt{-g}&\left[-\frac{1}{4}(1+\delta_{3})F^{\mu\nu}F_{\mu\nu}-\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu}\right.\\ &\left.-g^{\mu\nu}\left(\partial_{\mu}-ieA_{\mu}\right)\phi^{*}\left(\partial_{\nu}+ieA_{\nu}\right)\phi-(m_{\phi}^{2}+\xi R)\phi^{*}\phi\right].\end{split}$ (19)
@@
$\left\langle J^{\phi}_{z}\right\rangle_{\mathrm{reg}}=\sum_{i=0}^{3}(-1)^{i}\left\langle J_{z}\right\rangle_{i}=aH\frac{e^{2}E}{4\pi^{2}}\left[\frac{1}{6}\ln\frac{\Lambda^{2}}{H^{2}}-\frac{2\lambda^{2}}{15}+F_{\phi}(\lambda,\mu)\right]\,$, (29)
@@ 3.1 Behavior of the scalar current
and which depends only on $\lambda$, $\overline{m}_{\phi}$, and $\xi$. In Fig. 2 we show $\mathcal{J}_{\phi}$ for a minimally coupled scalar with $\xi=0$ (solid) and a conformally coupled scalar with $\xi=1/6$ (dashed) as a function of $\lambda$ (left) for various fixed values of $\overline{m}_{\phi}$ and as a function of $\overline{m}_{\phi}$ (right) for various fixed values of $\lambda$. We can then examine the behavior in various limits.
Figure 2: Left: Induced (dimensionless) Schwinger current defined in Eq. (31) for a minimally coupled scalar with $\xi=0$ (solid) and a conformally coupled scalar with $\xi=1/6$ (dashed) as a function of $\lambda\equiv eE/H^{2}$ for various values of fixed $\overline{m}_{\phi}\equiv m_{\phi}/H$ as labeled in the figure. Right: Same as left but now as a function of $\overline{m}_{\phi}$ for various fixed values of $\lambda$.
@@
where $\gamma_{E}\approx 0.577$ is Euler’s constant. The two terms in the first parenthesis come from the regularized current while the terms in the second parenthesis come from the counterterm. The first three terms in the second parenthesis arise due to the curvature corrections to the electron mass in the propagator. So we see explicitly the curvature corrections to the conformally coupled scalar mass entering in $\delta^{\phi}_{3}$ are crucial (otherwise the $\log$ and $\coth^{-1}$ terms cancel as $\overline{m}_{\phi}\to 0$) for compensating the negative terms and ensuring a positive renormalized current. We also see that in the massless limit the current for a conformally coupled scalar converges to a constant that is fixed by the physical renormalization in Eq. (14) and numerically equal to $\mathcal{J}_{\phi}\approx 0.011\lambda$. As we’ll see in the next section, for a conformally coupled scalar the physical current in Eq. (3) has the same behaviour as the charged fermion current (see Fig. 3), which gives us further confidence in our renormalization procedure.
@@ 3.1.3 Large mass limit
We now consider the case in which the charged scalar particles are very heavy in the limit $\overline{m}_{\phi}^{2}\gg(\lambda,12\xi)$. From Eq. (3) this gives for the renormalized current
@@
where the integral function is given by
@@
We have found that the result for the regularized current, once we take the large mass limit for the charged field, reproduces what we would obtain from the Euler-Heisenberg Lagrangian. Recall that computing $\langle J_{\mu}\rangle$ means computing the expectation value of an operator which in turn is defined in terms of creation and annihilation operators describing the quantum fluctuations of the scalar field. In the large mass limit these quantum fluctuations are too heavy to become real pairs of charged particles. However, they still contribute to the expectation value of the current operator and the result matches what one would obtain by first integrating out the heavy field and including the effects of its fluctuations in the effective Euler-Heisenberg Lagrangian. While the flat part of this effective Lagrangian has been long known to all orders, the curved part is difficult to derive and to our knowledge is known only to leading $\overline{m}_{\phi}^{-2}$ order Bastianelli:2008cu . Given that our method reproduces exactly this lowest order we speculate that it also reproduces correctly all the higher orders in $\overline{m}_{\phi}^{-2n}$ (with $n$ a positive integer), but leave a detailed investigation of this point to future work. The matching between $\left\langle J^{\phi}_{z}\right\rangle_{\mathrm{reg}}$ in the large mass limit and the Euler-Heisenberg effective currents was already pointed out in Banyeres:2018aax . Our derivation is subtly different in the way we separate the contribution of the counterterm which as we saw should not be included in the comparison and how we consistently keep track from the very beginning of the terms proportional to the $\xi$ coupling of the scalar field to the Ricci scalar.
@@ 4 Fermion Schwinger current in de Sitter
$\left\langle J^{\psi}_{z}\right\rangle_{\mathrm{reg}}=2aH\frac{e^{2}E}{4\pi^{2}}\left[\frac{1}{3}\ln\frac{\Lambda^{2}}{H^{2}}+\frac{1}{2}+\frac{2\lambda^{2}}{15}+\frac{3\overline{m}_{\psi}^{2}}{2\lambda^{2}}\left(1+\frac{x}{2\lambda}\ln\left(\frac{x-\lambda}{x+\lambda}\right)\right)+F_{\psi}(\lambda,\overline{m}_{\psi})\right].$~~~~~~ (59)
@@ 4.1 Behavior of the fermion current
In Fig. 3 we show $\mathcal{J}_{\psi}$ (solid) as a function of $\lambda$ (left) for various fixed values of $\overline{m}_{\psi}$ and as a function of $\overline{m}_{\psi}$ (right) for various fixed values of $\lambda$. For comparison we also show the conformally coupled scalar (dashed) case presented in the previous section. We see very similar behavior between the fermion and conformally coupled scalar currents.
Figure 3: Left:  The induced (dimensionless) Schwinger current defined in Eq. (61) for a charged fermion as a function of $\lambda\equiv eE/H^{2}$ for various fixed values of $\overline{m}_{\psi}\equiv m_{\psi}/H$ as shown. Right: Same as left, but now as a function of $\overline{m}_{\psi}$ for various fixed values of $\lambda$ as shown. For comparison we also show the conformally coupled scalar (dashed) case presented in the previous section.
@@ 4.1.3 Large mass limit
In the large mass limit $\overline{m}_{\psi}\gg 1$ we obtain,
@@ 5 Summary and conclusions
With this counterterm we then recomputed the induced Schwinger current in de Sitter for both charged fermions and minimally coupled scalars finding them to be finite and positive including in the small mass regime. In particular, our results are free of the peculiar negative IR divergences found in previous calculations Kobayashi:2014zza ; Hayashinaka:2016dnt ; Hayashinaka:2016qqn ; Banyeres:2018aax . We have traced the origin of these negative IR divergences to renormalization conditions which are inconsistent with the existence of a constant classical electric field in de Sitter space. We have also computed the Schwinger current in de Sitter space for a conformally coupled scalar again finding it to be positive (in contrast to Hayashinaka:2018amz ) and finite with very similar behavior to the charged fermion current. We then studied the behavior and various limits of the scalar and fermion Schwinger currents (see Fig. 2 and Fig. 3). We find the same IR hyper-conductivity for a light minimally coupled scalar found in previous calculations Kobayashi:2014zza ; Banyeres:2018aax and recover the result for the current in flat space in the strong field limit. However, in the weak field and small mass limits our results differ from previous calculations and in particular, are IR finite. Since the current is IR safe and always positive, we can consider production of massless (conformally coupled) scalars and fermions during inflation, something which cannot occur from purely gravitational production.
@@
Finally, we emphasize that since the gauge field is not quantized (only the charged fields), the one-loop vacuum polarization calculation to fix the counterterm contains the full quantum information so our results are non-perturbative. While the calculation of the regularized current is exact, the calculation of the counterterm is done with an approximation of the vacuum polarization diagram in de Sitter space, as discussed in Section 2.3. Thus our final result for the renormalized current is non-perturbative but not exact, and can in principle be improved with a full calculation of the counterterm in de Sitter space. Nevertheless, our results resolve the puzzling IR behavior found previously in the literature and are an important step in gaining understanding of the Schwinger effect in de Sitter space.
Acknowledgements.
The authors are especially thankful to Guillem Doménech, Jaume Garriga, and Takeshi Kobayashi for extensive discussions. The authors also thank Pedro Garcia Osorio, Manel Masip, and Jose Santiago for useful comments and discussions. This work has been partially supported by Junta de Andalucía Project A-FQM-472-UGR20 (fondos FEDER) and by SRA (10.13039/501100011033) and ERDF under grant PID2022-139466NB-C21 (R.V.M.) as well as by PID2022-140831NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER,UE (M.B.G.,A.T.M.), FCT CERN grant 10.54499/2024.00252.CERN (M.B.G.,A.T.M.,P.F), and FCT fellowship SFRH/BD/151475/2021 with DOI identifier 10.54499/SFRH/BD/151475/2021 (P.F.). LU is supported by the Slovenian Research Agency under the research core funding No.P1-0035, and by the research grants J1-60026 and J1-4389. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). RVM thanks the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and partial support during the completion of this work.
@@ Appendix B Calculation of the QED Counterterm ($\delta_{3}$)
Here we present details of the calculation of the counterterm obtained by computing the vacuum polarization diagram (see Fig. 1) for the photon 2-point function following Ch.16.2 of Schwartz:2014sze .
@@ B.1 Vacuum polarization in scalar QED
Finally, as $\Pi_{2}^{\mu\nu}=\left(-p^{2}g^{\mu\nu}\right)e^{2}\Pi_{2}(p^{2})$ and imposing the renormaliztion condition in Eq. (14) we find the counterterm to be (defining $\overline{M}_{\phi}\equiv M_{\phi}/H$),
$\delta_{3}^{\phi}$ $=$ $-e^{2}\Pi(2H^{2})$
$=$ $\left(\frac{e}{12\pi}\right)^{2}\Big{[}-3\ln(\frac{\Lambda^{2}}{H^{2}})+3\ln(\overline{M}_{\phi}^{2})+6\left(2\overline{M}_{\phi}^{2}+1\right)^{3/2}\coth^{-1}\left(\sqrt{2\overline{M}_{\phi}^{2}+1}\right)-12\overline{M}_{\phi}^{2}-8\Big{]}\,.$
Note that in the limit $\overline{M}_{\phi}\to 0$ there is no IR divergence and $\delta_{3}^{\phi}$ converges to,
$\delta_{3}^{\phi}$ $=$ $-e^{2}\Pi(2H^{2})=\left(\frac{e}{12\pi}\right)^{2}\Big{[}-3\ln(\frac{\Lambda^{2}}{H^{2}})+\ln 8-8\Big{]}\,.$

### v2

@@ 3 Scalar Schwinger current in de Sitter
$S$ $=$ $\int d^{4}x\sqrt{-g}\Big[-\frac{1}{4}(1+\delta_{3})F^{\mu\nu}F_{\mu\nu}-\frac{1}{2}m_{A}^{2}A_{\mu}A^{\mu}$ (27)
- $g^{\mu\nu}\left(\partial_{\mu}-ieA_{\mu}\right)\phi^{*}\left(\partial_{\nu}+ieA_{\nu}\right)\phi-(m_{\phi}^{2}+\xi R)\phi^{*}\phi\Big].$
@@
$\left\langle J^{\phi}_{z}\right\rangle_{\mathrm{reg}}=\sum_{i=0}^{3}(-1)^{i}\left\langle J_{z}\right\rangle_{i}=aH\frac{e^{2}E}{4\pi^{2}}\left[\frac{1}{6}\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)-\frac{2\lambda^{2}}{15}+F_{\phi}(\lambda,\mu)\right]\,$, (37)
@@ 3.1 Behavior of the scalar current
and which depends only on $\lambda$, $\overline{m}_{\phi}$, and $\xi$. In Fig. 2 we show $\mathcal{J}_{\phi}$ for a minimally coupled scalar with $\xi=0$ (solid) and a conformally coupled scalar with $\xi=1/6$ (dashed) as a function of $\lambda$ (top) for various fixed values of $\overline{m}_{\phi}$ and as a function of $\overline{m}_{\phi}$ (bottom) for various fixed values of $\lambda$. We can then examine the behavior in various limits.
Figure 2: Top: Induced (dimensionless) Schwinger current defined in Eq. (39) for a minimally coupled scalar with $\xi=0$ (solid) and a conformally coupled scalar with $\xi=1/6$ (dashed) as a function of $\lambda\equiv eE/H^{2}$ for various values of fixed $\overline{m}_{\phi}\equiv m_{\phi}/H$ as labeled in the figure. Bottom: Same as top but now as a function of $\overline{m}_{\phi}$ for various fixed values of $\lambda$.
@@
where $\gamma_{E}\approx 0.577$ is Euler’s constant. The two terms in the first parenthesis come from the regularized current while the terms in the second parenthesis come from the counterterm. The first three terms in the second parenthesis arise due to the curvature corrections to the electron mass in the propagator. So we see explicitly the curvature corrections to the conformally coupled scalar mass entering in $\delta^{\phi}_{3}$ are crucial (otherwise the $\log$ and $\coth^{-1}$ terms cancel as $\overline{m}_{\phi}\to 0$) for compensating the negative term and ensuring a positive renormalized current. We also see that in the massless limit the current for a conformally coupled scalar converges to a constant that is fixed by the physical renormalization in Eq. (11) and numerically equal to $\mathcal{J}_{\phi}\approx 0.011\lambda$. As we’ll see in the next section, for a conformally coupled scalar the physical current in Eq. (3) has the same behaviour as the charged fermion current (see Fig. 3), which gives us further confidence in our renormalization procedure.
@@ 3.1.3 Large mass limit
We now consider the case in which the charged scalar particles are very heavy. This limit was already largely discussed in Banyeres:2018aax , where it was pointed out that the resulting current matches the one obtained from the Euler-Heisenberg effective action. Although our final results are similar, our derivation has some differences compared to that of Banyeres:2018aax . First, we separate the contribution of the counterterm which should not be included in the comparison, as we explain below; second, we consistently keep track from the very beginning of the terms proportional to the $\xi$ coupling of the scalar field to the Ricci scalar. The latter point results in several terms with powers of $\xi$ which were not accounted for in Banyeres:2018aax .
Starting from Eq. (3) and taking the limit $\overline{m}_{\phi}^{2}\gg(\lambda,12\xi)$ gives
@@
where the integral function is given by,
@@
We have found that the result for the regularized current, once we take the large mass limit for the charged field, reproduces what we would obtain from the Euler-Heisenberg Lagrangian. Recall that computing $\langle J_{\mu}\rangle$ means computing the expectation value of an operator which in turn is defined in terms of creation and annihilation operators describing the quantum fluctuations of the scalar field. In the large mass limit these quantum fluctuations are too heavy to become real pairs of charged particles. However, they still contribute to the expectation value of the current operator and the result matches what one would obtain by first integrating out the heavy field and including the effects of its fluctuations in the effective Euler-Heisenberg Lagrangian. While the flat part of this effective Lagrangian has been long known to all orders, the curved part is difficult to derive and to our knowledge is known only to leading $\overline{m}_{\phi}^{-2}$ order Bastianelli:2008cu . Given that our method reproduces exactly this lowest order we speculate that it also reproduces correctly all the higher orders in $\overline{m}_{\phi}^{-2n}$ (with $n$ a positive integer), but leave a detailed investigation of this point to future work.
@@ 4 Fermion Schwinger current in de Sitter
$\left\langle J^{\psi}_{z}\right\rangle_{\mathrm{reg}}=2aH\frac{e^{2}E}{4\pi^{2}}\left[\frac{1}{3}\ln\Big(\frac{\Lambda^{2}}{H^{2}}\Big)+\frac{1}{2}+\frac{2\lambda^{2}}{15}+\frac{3\overline{m}_{\psi}^{2}}{2\lambda^{2}}\left(1+\frac{x}{2\lambda}\ln\left(\frac{x-\lambda}{x+\lambda}\right)\right)+F_{\psi}(\lambda,\overline{m}_{\psi})\right].$~~~~~~ (67)
@@ 4.1 Behavior of the fermion current
In Fig. 3 we show $\mathcal{J}_{\psi}$ (solid) as a function of $\lambda$ (top) for various fixed values of $\overline{m}_{\psi}$ and as a function of $\overline{m}_{\psi}$ (bottom) for various fixed values of $\lambda$. For comparison we also show the conformally coupled scalar (dashed) case presented in the previous section. We see very similar behavior between the fermion and conformally coupled scalar currents.
Figure 3:  Top: Induced (dimensionless) Schwinger current defined in Eq. (69) for a charged fermion as a function of $\lambda\equiv eE/H^{2}$ for various fixed values of $\overline{m}_{\psi}\equiv m_{\psi}/H$ as shown. Bottom: Same as top, but as a function of $\overline{m}_{\psi}$ for fixed values of $\lambda$ as shown. For comparison we also show the conformally coupled scalar (dashed) case presented in the previous section.
@@ 4.1.3 Large mass limit
The derivation in this section agrees with and confirms the results obtained in Appendix C of Banyeres:2018aax . In the large mass limit $\overline{m}_{\psi}\gg 1$ we obtain,
@@ 5 Summary and conclusions
With this counterterm we then recomputed the induced Schwinger current in de Sitter for both charged fermions and minimally coupled scalars finding them to be finite and positive including in the small mass regime. In particular, our results are free of the peculiar negative IR divergences found in previous calculations Kobayashi:2014zza ; Hayashinaka:2016dnt ; Hayashinaka:2016qqn . In Banyeres:2018aax the authors argued that the negative current is not physical and correctly connected the origin of the issue with the renormalization condition, but did not connect such a condition to the tachyonic mass of the vector, which is a crucial point of our work. Here we have examined this issue in detail and traced the origin of these negative IR divergences to renormalization conditions which are inconsistent with the existence of a constant classical electric field in de Sitter space. We have also computed the Schwinger current in de Sitter space for a conformally coupled scalar again finding it to be positive (in contrast to Hayashinaka:2018amz ) and finite with very similar behavior to the charged fermion current. We then studied the behavior and various limits of the scalar and fermion Schwinger currents (see Fig. 2 and Fig. 3). We find the same IR hyper-conductivity for a light minimally coupled scalar found in previous calculations Kobayashi:2014zza ; Banyeres:2018aax and recover the result for the current in flat space in the strong field limit. However, in the weak field and small mass limits our results differ from previous calculations and in particular, are IR finite. Since the current is IR safe and always positive, we can consider production of massless (conformally coupled) scalars and fermions during inflation, something which cannot occur from purely gravitational production.
@@
Finally, we emphasize that since the gauge field is not quantized (only the charged fields), the one-loop vacuum polarization calculation to fix the counterterm contains the full quantum information so our results are non-perturbative. While the calculation of the regularized current is exact, the calculation of the counterterm is done with an approximation of the vacuum polarization diagram in de Sitter space, as discussed in Section 2.3. Thus our final result for the renormalized current is non-perturbative but not exact and can in principle be improved with a full calculation of the counterterm in de Sitter space. Nevertheless, our results resolve the puzzling IR behavior found previously in the literature and are an important step in gaining understanding of the Schwinger effect in de Sitter space.
Acknowledgments:  The authors are especially thankful to Guillem Doménech, Jaume Garriga, and Takeshi Kobayashi for extensive discussions. The authors also thank Pedro Garcia Osorio, Manel Masip, and Jose Santiago for useful comments and discussions. This work has been partially supported by Junta de Andalucía Projects  P21-00199, A-FQM-472-UGR20 (fondos FEDER) and by SRA (10.13039/501100011033) and ERDF under grant PID2022-139466NB-C21 (R.V.M.) as well as by PID2022-140831NB-I00 funded by MICIU/AEI/10.13039/501100011033 and FEDER,UE (M.B.G.,A.T.M.), FCT CERN grant 10.54499/2024.00252.CERN (M.B.G.,A.T.M.,P.F), and FCT fellowship SFRH/BD/151475/ 2021 with DOI identifier 10.54499/SFRH/BD/151475/2021 (P.F.). LU is supported by the Slovenian Research Agency under the research core funding No.P1-0035, and by the research grants J1-60026 and J1-4389. This article is based upon work from COST Action COSMIC WISPers CA21106, supported by COST (European Cooperation in Science and Technology). RVM thanks the Mainz Institute for Theoretical Physics (MITP) of the Cluster of Excellence PRISMA+ (Project ID 390831469), for its hospitality and partial support during the completion of this work.
@@ Appendix B Calculation of the QED Counterterm ($\delta_{3}$)
Here we present details of the calculation of $\delta_{3}$ obtained by computing the vacuum polarization diagram (see Fig. 1) for the photon 2-point function following Ch.16.2 of Schwartz:2014sze .